# Effects of orbital eccentricity on continuous gravitational waveforms from triaxially-deformed precessing neutron stars in tight binaries

### v1

@@ Abstract
The successful detection of continuous gravitational waves (GWs) from spinning neutron stars (NSs) will shape our understanding of the physical properties of dense matter under extreme conditions. Binary population synthesis simulations show that forthcoming space-borne GW detectors may be capable of detecting some tight Galactic double NSs (DNSs) with 10-minute orbital periods. Successfully searching for continuous GWs from such a close DNS demands extremely precise waveform templates considering the interaction between the NS and its companion. Unlike the isolated formation channel, the DNSs from the dynamical formation channel have moderate to high orbital eccentricities. To accommodate these systems, we generalize the analytical waveforms from triaxial nonaligned NSs under spin-orbit coupling derived by Feng et al. [Phys. Rev. D 108, 063035 (2023)] to incorporate the effects of the orbital eccentricity. Our findings suggest that for DNSs formed through isolated binary evolution, the impact of eccentricity on the continuous GWs of their NSs can be neglected. In contrast, for DNSs formed through dynamical processes, it is necessary to consider eccentricity, as high-eccentricity orbits can result in a fitting factor of $\lesssim 0.97$ within approximately 0.5 to 2 years of a coherent search. Once the GWs from spinning NSs in tight binaries are detected, the relative measurement accuracy of eccentricity can reach $\Delta e/e\sim O(10^{-7})$ for a signal-to-noise ratio of $O(100)$ based on the Fisher information matrix, bearing significant implications for understanding the formation mechanisms of DNSs.
@@ I Introduction
The current continuous waveform templates used in the LIGO/Virgo/KAGRA Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2020) searches are predominantly based on isolated NSs Abbott et al. (2022, 2021a, 2022a, 2022b) and NSs in binary systems with long orbital periods, where interactions with the companion star are neglected Abbott et al. (2021b); Covas and Sintes (2020, 2019); Zhang et al. (2021); Leaci and Prix (2015); Pagliaro et al. (2023); Covas et al. (2024). At present, there is still no credible detection of continuous GWs. However, the upper limits on GW emission from various sources and sky regions are constrained Ming et al. (2024); Owen et al. (2024); Liu and Zou (2022); Rajbhandari et al. (2021); Dergachev et al. (2019); Zhu et al. (2016).
Binary population synthesis simulations based on the classical isolated binary evolution predict that upcoming space-based GW detectors, such as LISA Amaro-Seoane et al. (2017), TianQin Luo et al. (2016), and Taiji Hu and Wu (2017), have a promising capability to detect binaries with orbital period $P_{\rm b}\sim 10~{}{\rm min}$ and eccentricity $e\sim 0.01$ Andrews et al. (2020); Miao et al. (2021); Wagg et al. (2022); Feng et al. (2023a). The next-generation ground-based GW detectors, such as Einstein Telescope Punturo et al. (2010); Kalogera et al. (2021) and Cosmic Explorer Srivastava et al. (2022), which will operate almost concurrently with LISA, TianQin, and Taiji, are very likely to detect spinning NSs in these binary systems. The continuous waveforms of the spinning NS components in these host binaries are shaped by various factors, one of which is the spin-orbit coupling Poisson and Will (2014). Binaries with spins not aligned with the orbital angular momentum will precess around the direction of the total angular momentum, thereby introducing complexities in waveform patterns Feng et al. (2023b). Accurate modeling of these patterns is crucial for detecting and characterizing the GWs emitted by such systems Feng et al. (2024). Due to their low orbital eccentricities, the waveform modeling of spinning NSs in DNS systems formed in isolation within the Galactic field, as per Ref. Feng et al. (2023b), has been calculated for circular orbits during spin precession.
@@ II Gravitational waveforms from a triaxial NS under spin precession
Considering a DNS system with component masses $m_{1,2}$, and an eccentric orbit characterized by a semimajor axis $a$ and eccentricity $e$. For the simple precession scenario with secular evolution Apostolatos et al. (1994); Barker and O’Connell (1975), where only one of the DNS has spin angular momentum $\boldsymbol{S}$ and the spin-spin interaction is neglected Poisson and Will (2014); Feng et al. (2023b), the precession equation for $\boldsymbol{S}$ takes a simplified form after averaging over one eccentric orbit:
@@
Figure 1: The geometry of the $J$-aligned frame and $S$-aligned frame. Both are fixed to the spinning NS with $O_{1}$ as the coordinate origin Feng et al. (2023b). The opening angle for the precession cone of $\boldsymbol{S}$ is denoted by $\theta_{S}$. The $X_{S}$ axis is set to initially coincide with the $X_{J}$ axis, and $\boldsymbol{S}(t=0)$ is within the $Y_{J}-Z_{J}$ plane. The projection of $\boldsymbol{S}(t)$ onto the $X_{J}-Y_{J}$ plane is along ${O_{1}N}$, and it precesses with an angle $\alpha(t)$, given in Eq. (3). The distant observer, represented by $\boldsymbol{D}$, is positioned at colatitude $\iota$ from $\boldsymbol{J}$ within the $Y_{J}-Z_{J}$ plane.
Under the transverse-traceless gauge, the metric perturbation can be expressed as a combination of two GW polarizations, given by $h_{jk}^{\mathrm{TT}}=h_{+}\left(\hat{e}_{+}\right)_{jk}+h_{\times}\left(\hat{e}_{\times}\right)_{jk}$. Here, the polarization tensors are $\hat{e}_{+}\equiv\hat{v}\otimes\hat{v}-\hat{w}\otimes\hat{w},\,\hat{e}_{\times}\equiv\hat{v}\otimes\hat{w}+\hat{w}\otimes\hat{v}$, with $\hat{v}\equiv\hat{e}_{y}\cos\iota-\hat{e}_{z}\sin\iota$ and $\hat{w}\equiv-\hat{e}_{x}$ forming the orthogonal basis vectors transverse to the direction of wave propagation. The set $(\hat{e}_{x},\hat{e}_{y},\hat{e}_{z})$ serves as the basis vectors for the J-aligned frame. The formulations for the two GW polarizations are elaborated in Ref. Zimmermann (1980),
@@
where Einstein summation is applied to the indices $\mu$ and $\nu$, each ranging over the set $\{1,2,3\}$. The matrix $\mathcal{R}$ denotes the rotation matrix that transforms from the body frame (constructed by three principal axes of inertia) of the spinning NS to the $J$-aligned inertial frame. This matrix is achieved through the following rotation transformations: initially from the body frame to the $S$-aligned frame using the matrix $R$ (see Eq. (24) in Ref. Zimmermann (1980)), followed by a transformation from the $S$-aligned frame to the $J$-aligned frame using $\mathcal{T}_{S\rightarrow J}$. Therefore,
@@
The gravitational waveforms emitted by the spinning NS under spin precession are closely related to the spin angular velocity, making it crucial to solve Eq. (1). Assume that the spinning NS is a rigid body, with a spin period of $P_{\rm s}$, and the magnetic dipole moment is $p$ (in the Gaussian CGS unit). Other possible sources of external torque that can be added to the right-hand side of Eq. (1) include gravitational radiation and magnetic dipole radiation, whose magnitudes are given by the following expressions Ostriker and Gunn (1969); Gao et al. (2023):
@@
For a spinning NS with a magnetic field of $10^{9}~{}\rm{G}$ and a spin period of $10~{}\rm{ms}$ in a tight DNS with a 10-minute orbital period (see Table 1), the rates for the change of angular velocity are
@@
Waveforms from a triaxial NS are typically represented through a series expansion using certain small parameters, as detailed in Refs. Zimmermann (1980); Van Den Broeck (2005); Gao et al. (2020); Feng et al. (2023b). To simplify the calculations, we specify the angular frequencies of free precession $\Omega_{\rm p}$ and rotation $\Omega_{\rm r}$ for the spinning NS, along with three small parameters (oblateness or poloidal ellipticity $\epsilon$, non-sphericity $\kappa$, and wobble angle $\gamma$) delineating the NS’s characteristics
@@
Figure 2: The amplitude envelope (taking only positive values) of $h_{+}^{(1)}$ for different eccentricities, assuming a binary orbital period of 10 minutes, with the NS parameters listed in Table 1.
@@
Non-sphericity $\kappa$ $1.75\times 10^{-4}$
@@
III Comparison with waveforms from circular-orbit case
@@
[…]
@@
Figure 3: The time required for the fitting factor (FF) to decrease to 0.97 for the first three NS waveform components in binaries [see Eqs. (14)]. The parameters used here for the spinning NS can be found in Table 1. The red solid lines in each subplot correspond to contours for 0.2, 0.5, 1, 5, and 10 years. The abscissa and ordinate are the orbital period ($P_{\rm b}$) and eccentricity ($e$) of the binary respectively, and the top abscissa is the second harmonic frequency of orbit ($f_{\rm g}=2/P_{\rm b}$). The ranges of $P_{\rm b}\in[5,120]~{}{\rm min}$ and $e\in[0.01,0.8]$ fall in the mHz band of the space-borne GW detectors. The blue contours represent the DNS population from COMPAS simulation through isolated binary evolution Wagg et al. (2022), while the orange and magenta contours denote two putative fast-merging formation channels, i.e., fast-merging isolated binaries and dynamically formed binaries, respectively Andrews et al. (2020). These three types of contours indicate the percentage of the enclosed DNS population, from inner to outer, ranging from $10\%$ to $80\%$ in $10\%$ increments. The black dashed line in the subplot illustrates the evolution track of the Hulse-Taylor binary (PSR B1913+16), assuming that orbital decay results exclusively from gravitational radiation. For a tight binary system with a 10-minute orbital period and an eccentricity below 0.1, the waveforms emitted by the component NS in an eccentric binary remain closely aligned with those from a circular orbit for 5 years of observation.
@@
Figure 3 illustrates the time it takes for the FF to decrease to 0.97 for the first three NS waveform components in Eqs. (14). The horizontal and vertical axes represent the orbital periods ($P_{\rm b}$) and eccentricities ($e$) of the binaries, respectively. The secondary horizontal axis on top denotes the second harmonic frequency of orbital motion ($f_{\rm g}=2/P_{\rm b}$). These panels present DNS populations with equal component masses, $m_{1}=m_{2}=1.4M_{\odot}$. The orbital period of these binaries ranges from 5 minutes to 2 hours, covering the frequency band of space-based GW detectors (e.g., LISA, TianQin, and Taiji), with eccentricities taken from 0.01 up to 0.8. The characteristic parameters of the spinning NS are given in Table 1 Feng et al. (2023b). In each subplot, the red solid lines indicate contours for 0.2, 0.5, 1, 5, and 10 years. As seen in this figure, in the case of a tight binary system with a 10-minute orbital period and an eccentricity less than 0.1, the waveforms emitted by the NS components remain closely aligned for 5 years, with no significant difference from the circular orbit. The subfigures also depict the $P_{\rm b}-e$ evolution track of the Hulse-Taylor binary (PSR B1913+16) Hulse and Taylor (1975) within the parameter space as a black dashed line. The results indicate that we do not need to consider the impact of residual eccentricity on the GW radiation waveform of NSs in Hulse-Taylor-like binaries when they enter the relevant frequency band (see e.g., Ref. Zhao et al. (2021)).
@@
Figure 3 also depicts the parameter distributions in $P_{b}-e$ space provided by three DNS population models. These three types of contours indicate the percentage of the enclosed DNS population, from inner to outer, ranging from $10\%$ to $80\%$ in $10\%$ increments. The first model, detectable by LISA and represented by the blue contours, originates from Ref. Wagg et al. (2022). This model (abbreviated as COMPAS model) simulates compact binary populations in the Milky Way using the rapid population synthesis code COMPAS Riley et al. (2022), based on the classical isolated binary evolution channel. The simulation incorporates a novel, empirically-based analytical model for the metallicity-dependent star formation history of our Galaxy. Although most of the DNSs in the simulation have an equal mass of approximately $1.3~{}M_{\odot}$, it is expected that this should not significantly deviate from the $1.4~{}M_{\odot}$ case that we adopt in this study. For this DNS population model, we find that for over $80\%$ of the sources (where $e<0.3$), a noticeable deviation from the circular orbit case would only occur after more than 10 years of observation. The second model (denoted by orange contours) and the third model (denoted by magenta contours) come from the putative fast-merging DNS evolutionary channel Andrews et al. (2020). For these two models, DNSs are randomly created by using a log-normal distribution for orbital separation with parameters $\mu=-0.1$ and $\sigma=0.2$ in units of Solar radius. The eccentricity follows a normal distribution with means of $\mu=0.2$ and $\mu=0.7$ for the two models, both having a standard deviation of $\sigma=0.1$ and truncated to ensure that $10<e<$ 1. For the second model with fast-merging isolated binaries (abbreviated as IsolatedFast model), the majority of the DNSs with eccentricities less than 0.3 do not exhibit deviations from the circular orbit within 10 years of observation. For the third model, which involves fast-merging dynamically formed binaries (abbreviated as DynamicalFast model), most sources will exhibit deviations from the circular orbit case within 0.5 to 2 years for $h_{+,\times}^{(1)}$. For $h_{+,\times}^{(2)}$ and $h_{+,\times}^{(3)}$, deviations are expected within 1 to 5 years.

### v2

# Effects of orbital eccentricity on continuous gravitational waveforms from triaxially deformed precessing neutron stars in tight binaries

@@ Abstract
The successful detection of continuous gravitational waves from spinning neutron stars (NSs) will shape our understanding of the physical properties of dense matter under extreme conditions. Binary population synthesis simulations show that forthcoming space-borne gravitational wave detectors may be capable of detecting some tight Galactic double NSs with 10-min orbital periods. Successfully searching for continuous waves from the individual NS in such a close binary demands extremely precise waveform templates considering the interaction between the NS and its companion. Unlike the isolated formation channel, double NS systems from the dynamical formation channel have moderate to high orbital eccentricities. To accommodate these systems, we generalize the analytical waveforms from triaxial nonaligned NSs under spin-orbit coupling derived by Feng et al. [Phys. Rev. D 108, 063035 (2023)] to incorporate the effects of the orbital eccentricity. Our findings suggest that for binaries formed through isolated binary evolution, the impact of eccentricity on the continuous waves of their NSs can be neglected. In contrast, for those formed through dynamical processes, it is necessary to consider eccentricity, as high-eccentricity orbits can result in a fitting factor of $\lesssim 0.97$ (0.9) within approximately 0.5 (1) to 2 (5) yr of a coherent search (at wave frequencies of 100 and 200 Hz). Once the continuous waves from spinning NSs in tight binaries are detected, the relative measurement accuracy of eccentricity can reach $\Delta e/e\sim O(10^{-7})$ for a signal-to-noise ratio of $O(100)$ based on the Fisher information matrix, bearing significant implications for understanding the formation mechanisms of double NS systems.
@@ I Introduction
The current continuous waveform templates of NSs used in the LIGO/Virgo/KAGRA Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2020) searches are predominantly based on isolated NSs Abbott et al. (2022, 2021a, 2022a, 2022b) and NSs in binary systems with long orbital periods, where interactions with the companion star are neglected Abbott et al. (2021b); Covas and Sintes (2020, 2019); Zhang et al. (2021); Leaci and Prix (2015); Pagliaro et al. (2023); Covas et al. (2024). At present, there is still no credible detection of continuous GWs of NSs. However, the upper limits on GW emission from various sources and sky regions are constrained Ming et al. (2024); Owen et al. (2024); Liu and Zou (2022); Rajbhandari et al. (2021); Dergachev et al. (2019); Zhu et al. (2016).
Binary population synthesis simulations based on the classical isolated binary evolution predict that upcoming space-based GW detectors, such as LISA Amaro-Seoane et al. (2017), TianQin Luo et al. (2016), and Taiji Hu and Wu (2017), have a promising capability to detect binaries with orbital period $P_{\rm b}\sim 10~{}{\rm min}$ and eccentricity $e\sim 0.01$ Andrews et al. (2020); Miao et al. (2021); Wagg et al. (2022); Feng et al. (2023a). The next-generation ground-based GW detectors, such as Einstein Telescope Punturo et al. (2010); Kalogera et al. (2021) and Cosmic Explorer Srivastava et al. (2022), which will operate almost concurrently with LISA, TianQin, and Taiji, are very likely to detect spinning NSs in these binary systems. The continuous waveforms of the spinning NS components in these host binaries are shaped by various factors, one of which is the spin-orbit coupling Poisson and Will (2014). Binaries with spins not aligned with the orbital angular momentum will precess around the direction of the total angular momentum, thereby introducing complexities in waveform patterns Feng et al. (2023b). Accurate modeling of these patterns is crucial for detecting and characterizing the GWs emitted by such systems Feng et al. (2024). Because of their low orbital eccentricities, the waveform modeling of spinning NSs in double NS (DNS) systems formed in isolation within the Galactic field, as per Ref. Feng et al. (2023b), has been calculated for circular orbits during spin precession.
@@ II Gravitational waveforms from a triaxial NS under spin precession
Consider a DNS system with component masses $m_{1,2}$, and an eccentric orbit characterized by a semimajor axis $a$ and eccentricity $e$. For the simple precession scenario with secular evolution Apostolatos et al. (1994); Barker and O’Connell (1975), where only one component NS of the DNS has spin angular momentum $\boldsymbol{S}$ and the spin-spin interaction is neglected Poisson and Will (2014); Feng et al. (2023b), the precession equation for $\boldsymbol{S}$ takes a simplified form after averaging over one eccentric orbit:
@@
Figure 1: The geometry of the $J$- and $S$-aligned frames. Both are fixed to the spinning NS with $O_{1}$ as the coordinate origin Feng et al. (2023b). The opening angle for the precession cone of $\boldsymbol{S}$ is denoted by $\theta_{S}$. The $X_{S}$ axis is set to initially coincide with the $X_{J}$ axis, and $\boldsymbol{S}(t=0)$ is within the $Y_{J}-Z_{J}$ plane. The projection of $\boldsymbol{S}(t)$ onto the $X_{J}-Y_{J}$ plane is along ${O_{1}N}$, and it precesses with an angle $\alpha(t)$, given in Eq. (3). The distant observer, represented by $\boldsymbol{D}$, is positioned at colatitude $\iota$ from $\boldsymbol{J}$ within the $Y_{J}-Z_{J}$ plane.
Under the transverse-traceless gauge, the metric perturbation can be expressed as a combination of two GW polarizations, given by $h_{jk}^{\mathrm{TT}}=h_{+}\left(\hat{e}_{+}\right)_{jk}+h_{\times}\left(\hat{e}_{\times}\right)_{jk}$. Here, the polarization tensors are $\hat{e}_{+}\equiv\hat{v}\otimes\hat{v}-\hat{w}\otimes\hat{w},\,\hat{e}_{\times}\equiv\hat{v}\otimes\hat{w}+\hat{w}\otimes\hat{v}$, with $\hat{v}\equiv\hat{e}_{y}\cos\iota-\hat{e}_{z}\sin\iota$ and $\hat{w}\equiv-\hat{e}_{x}$ forming the orthogonal basis vectors transverse to the direction of wave propagation. The set $(\hat{e}_{x},\hat{e}_{y},\hat{e}_{z})$ serves as the basis vectors for the $J$-aligned frame. The formulations for the two GW polarizations from the spinning NS are elaborated in Ref. Zimmermann (1980),
@@
where Einstein summation is applied to the indices $\mu$ and $\nu$, each ranging over the set $\{1,2,3\}$. The matrix $\mathcal{R}$ denotes the rotation matrix that transforms from the body frame (constructed by three principal axes of inertia) of the spinning NS to the $J$-aligned inertial frame. This matrix is achieved through the following rotation transformations: initially from the body frame to the $S$-aligned frame using the matrix $R$ [see Eq. (24) in Ref. Zimmermann (1980)], followed by a transformation from the $S$-aligned frame to the $J$-aligned frame using $\mathcal{T}_{S\rightarrow J}$. Therefore,
@@
The gravitational waveforms emitted by the spinning NS under spin precession are closely related to the spin angular velocity, making it crucial to solve Eq. (1). Assume that the spinning NS is a rigid body, with a spin period of $P_{\rm s}$, and the magnetic dipole moment is $p$ (in the Gaussian centimeter-gram-second unit). Other possible sources of external torque that can be added to the right-hand side of Eq. (1) include gravitational radiation and magnetic dipole radiation, whose magnitudes are given by the following expressions Ostriker and Gunn (1969); Gao et al. (2023):
@@
For a spinning NS with a magnetic field of $10^{9}~{}\rm{G}$ and a spin period of $10~{}\rm{ms}$ in a tight DNS with a 10-min orbital period (see Table 1), the rates for the change of angular velocity are
@@
Waveforms from a triaxial NS are typically represented through a series expansion using certain small parameters, as detailed in Refs. Zimmermann (1980); Van Den Broeck (2005); Gao et al. (2020); Feng et al. (2023b). To simplify the calculations, we specify the angular frequencies of free precession $\Omega_{\rm p}$ and rotation $\Omega_{\rm r}$ for the spinning NS, along with three small parameters (oblateness or poloidal ellipticity $\epsilon$, nonsphericity $\kappa$, and wobble angle $\gamma$) delineating the NS’s characteristics
@@
Figure 2: The amplitude envelope (taking only positive values) of $h_{+}^{(1)}$ for different eccentricities, assuming a binary orbital period of 10 min, with the NS parameters listed in Table 1.
@@
Nonsphericity $\kappa$ $1.75\times 10^{-4}$
@@
III Comparison with waveforms from NS in circular-orbit case
@@
where ${\boldsymbol{\lambda}}$ represents the set of parameters characterizing the waveforms. For a quasimonochromatic signal, the inner product $\left(h_{\rm e},h_{\rm c}\right)$ simplifies to $\left(h_{\rm e},h_{\rm c}\right)\equiv\int_{0}^{T_{\rm obs}}h_{\rm e}(t)h_{\rm c}(t)\mathrm{d}t$, where $T_{\rm obs}$ denotes the observation time. We assume that the host binary, where the spinning NS is located, can be detected by the space-based GW detectors (e.g., LISA, TianQin, and Taiji). We choose a threshold of 0.97 (0.9) for FF, corresponding to a $\sim 10\%$ ($27\%$) reduction in the detection rate Apostolatos (1995); Mukherjee et al. (2023).
@@
Figure 3: The time required for the FF to decrease to 0.97 for the first three NS waveform components in binaries [see Eqs. (14)]. The parameters used here for the spinning NS can be found in Table 1. The red solid lines in each subplot correspond to contours for 0.2, 0.5, 1, 5, and 10 yr. The abscissa and ordinate are the orbital period ($P_{\rm b}$) and eccentricity ($e$) of the binary respectively, and the top abscissa is the second harmonic frequency of orbit ($f_{\rm g}=2/P_{\rm b}$). The ranges of $P_{\rm b}\in[5,120]~{}{\rm min}$ and $e\in[0.01,0.8]$ fall in the millihertz band of the space-borne GW detectors. The blue contours represent the DNS population from COMPAS simulation through isolated binary evolution Wagg et al. (2022), while the orange and magenta contours denote two putative fast-merging formation channels, i.e., fast-merging isolated binaries and dynamically formed binaries, respectively Andrews et al. (2020). These three types of contours indicate the percentage of the enclosed DNS population, from inner to outer, ranging from $10\%$ to $80\%$ in $10\%$ increments. The black dashed line in the subplot illustrates the evolution track of the Hulse-Taylor binary (PSR B1913+16), assuming that orbital decay results exclusively from gravitational radiation. For a tight binary system with a 10-min orbital period and an eccentricity below 0.1, the waveforms emitted by the component NS in an eccentric binary remain closely aligned with those from a circular orbit for 5 yr of observation.
@@
Figure 3 illustrates the time it takes for the FF to decrease to 0.97 for the first three NS waveform components in Eqs. (14). The horizontal and vertical axes represent the orbital periods ($P_{\rm b}$) and eccentricities ($e$) of the binaries, respectively. The secondary horizontal axis on top denotes the second harmonic frequency of orbital motion ($f_{\rm g}=2/P_{\rm b}$). These panels present DNS populations with equal component masses, $m_{1}=m_{2}=1.4M_{\odot}$. The orbital period of these binaries ranges from 5 min to 2 h, covering the frequency band of space-based GW detectors (e.g., LISA, TianQin, and Taiji), with eccentricities taken from 0.01 up to 0.8. The characteristic parameters of the spinning NS are given in Table 1 Feng et al. (2023b). In each subplot, the red solid lines indicate contours for 0.2, 0.5, 1, 5, and 10 yr. As seen in this figure, in the case of a tight binary system with a 10-min orbital period and an eccentricity less than 0.1, the waveforms emitted by the NS components remain closely aligned for 5 yr, with no significant difference from the circular orbit. The subfigures also depict the $P_{\rm b}-e$ evolution track of the Hulse-Taylor binary (PSR B1913+16) Hulse and Taylor (1975) within the parameter space as a black dashed line. The results indicate that we do not need to consider the impact of residual eccentricity on the GW radiation waveform of NSs in Hulse-Taylor-like binaries when they enter the relevant frequency band (see e.g., Ref. Zhao et al. (2021)).
@@
Figure 3 also depicts the parameter distributions in $P_{b}-e$ space provided by three DNS population models. These three types of contours indicate the percentage of the enclosed DNS population, from inner to outer, ranging from $10\%$ to $80\%$ in $10\%$ increments. The first model, detectable by LISA and represented by the blue contours, originates from Ref. Wagg et al. (2022). This model (abbreviated as COMPAS model) simulates compact binary populations in the Milky Way using the rapid population synthesis code COMPAS Riley et al. (2022), based on the classical isolated binary evolution channel. The simulation incorporates a novel, empirically based analytical model for the metallicity-dependent star formation history of our Galaxy. Although most of the DNSs in the simulation have an equal mass of approximately $1.3~{}M_{\odot}$, it is expected that this should not significantly deviate from the $1.4~{}M_{\odot}$ case that we adopt in this study. For this DNS population model, we find that for over $80\%$ of the sources (where $e<0.3$), a noticeable deviation from the circular-orbit case would only occur after more than 10 yr of observation. The second model (denoted by orange contours) and the third model (denoted by magenta contours) come from the putative fast-merging DNS evolutionary channel Andrews et al. (2020). For these two models, DNSs are randomly created by using a log-normal distribution for orbital separation with parameters $\mu=-0.1$ and $\sigma=0.2$ in units of solar radius. The eccentricity follows a normal distribution with means of $\mu=0.2$ and $\mu=0.7$ for the two models, both having a standard deviation of $\sigma=0.1$ and truncated to ensure that $10<e<$ 1. For the second model with fast-merging isolated binaries (abbreviated as IsolatedFast model), the majority of the DNSs with eccentricities less than 0.3 do not exhibit deviations from the circular orbit within 10 yr of observation. For the third model, which involves fast-merging dynamically formed binaries (abbreviated as DynamicalFast model), most sources will exhibit deviations from the circular-orbit case within 0.5-2 yr for $h_{+,\times}^{(1)}$. For $h_{+,\times}^{(2)}$ and $h_{+,\times}^{(3)}$, deviations are expected within 1-5 yr. For comparison, we have calculated the time required for FF to decline to 0.9 for various NS waveform components. In the DynamicalFast model, the majority of sources will display deviations from the circular-orbit case within 1-5 yr for $h_{+,\times}^{(1)}$. For $h_{+,\times}^{(2)}$ and $h_{+,\times}^{(3)}$, deviations are anticipated within 2-10 yr.